# Back to the Basics: Rethinking Issue-Commit Linking with LLM-Assisted Retrieval

### v1

@@ Abstract.
Inspired by these observations, we propose EasyLink, which utilizes a vector database as a modern Information Retrieval technique. To address the long-standing problem of the semantic gap between issues and commits, EasyLink leverages a large language model to rerank the commits retrieved from the database. Under our evaluation, EasyLink achieves an average Precision@1 of 75.91%, improving over the state-of-the-art by over four times. Additionally, this paper provides practical guidelines for advancing research in issue-commit link recovery.
@@
Jieke Shi is the corresponding author.
copyright: acmlicensedjournalyear: 2026conference: IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Prior studies (Bachmann and Bernstein, 2009; Romo et al., 2014) have revealed that many issue-commit links can be missing during the development of large-scale projects. Manually recovering these links is not only time-consuming but error-prone, even for experienced developers (Ruan et al., 2019). To address this, several studies (Zhang et al., 2023b; Lin et al., 2021; Dong et al., 2022; Ruan et al., 2019; Sun et al., 2017a; Rath et al., 2018) have proposed learning-based approaches to automatically recover issue-commit links, achieving strong performance on datasets collected from open-source repositories. The evaluation method requires the tool to distinguish fix commits (i.e., “true links”) from plausible but non-fix commits (i.e., “false links”). However, prior studies’ evaluations often lack realism. ❶ Some studies use an unrealistic time window to select commits. For example, studies (Ruan et al., 2019; Nguyen et al., 2012; Sun et al., 2017b; Bachmann et al., 2010) rely on a narrow 7-day time window to select potential fix commits, which may miss many true links. Another example involves studies (Dong et al., 2022; Rath et al., 2018) that require the issue close time to select commits, but in practice, when commits are missing, the issue may not have a close time. ❷ Another limitation is the unrealistic false link distributions in prior evaluation datasets. Some (Lin et al., 2021; Ruan et al., 2019) use evaluation datasets where the number of false and true links is equal, an unrealistic assumption since, in reality, false links far outnumber true links (Dong et al., 2022). Recently, Zhang et al. (Zhang et al., 2023b) addressed this issue by using an imbalanced dataset with a fixed number of false links per issue. Still, this method overlooks a crucial factor: a higher commit frequency results in a larger pool of plausible commits for the tool to differentiate from the actual fix commit, making the task inherently more difficult.
@@
In this paper, we propose EasyLink, a novel method that integrates modern IR techniques and Large Language Models (LLMs) to enhance issue-commit linking performance in realistic settings by capturing deeper semantic relationships between issues and commits. EasyLink operates in two stages. The first stage fetches a set of commits that share similarities to a given issue, and the second stage reranks them by their relevance to the issue. Concretely, in the first stage, EasyLink leverages recent advances in Information Retrieval by using a modern vector database (optimized for high-dimensional similarity search) (Douze et al., 2024) to fetch the most similar commits for each issue. In the second stage, EasyLink prompts an LLM (GPT-4o (OpenAI, 2024)) to rerank the results. We use LLMs for their strong ability to capture the semantic relationship between issues and commits, bridging the semantic gap, as the most similar commit does not guarantee that it is the fix for the issue (Zhang et al., 2023b; Lin et al., 2021; Dong et al., 2022). EasyLink achieves an average Precision@1 of 75.91%, outperforming EALink (Zhang et al., 2023b) by 4$\times$ in our realistic datasets and by 30% in Precision@1 on EALink’s original evaluation setup. We also conduct an ablation analysis on each of EasyLink’s two stages. Our results demonstrate that advances in IR, such as context-aware dense embeddings (Wang et al., 2020; Song et al., 2020; Neelakantan et al., 2022) and efficient vector search (Douze et al., 2024; Wang et al., 2021), have been largely overlooked in the software traceability literature. Notably, even the out-of-the-box use of a vector database achieves a high average Precision@1 of 64.03%. Additionally, the strong performance of the reranking step, which improves Precision@1 by 11%, highlights the capability of LLMs to bridge the semantic gap between issues and commits (Rath et al., 2018; Guo et al., 2017). Finally, we discuss the lessons learned from our work, such as the need to consider modern IR baselines, for future research on software traceability. Our implementation, along with the evaluation dataset, has been anonymized and made available at (Package, 2025).
@@
    A new state-of-the-art for issue-commit linking: We propose EasyLink, which leverages modern IR techniques, including an off-the-shelf database FAISS (Douze et al., 2024), and addresses the problem of semantic gap (Rath et al., 2018; Guo et al., 2017) by prompting an LLM to rerank the retrieved results. In our evaluation, EasyLink improves over EALink in average Precision@1 from 14% to 76%.
The rest of the paper is organized as follows. Section 2 outlines the background of this work, including the limitations of existing work. Section 3 details the evaluation dataset construction under the Realistic Distribution Setting (RDS). Section 4 introduces EasyLinkSection 5 describes the experimental setup. Section 6 presents the experimental results. Section 7 discusses the lessons learned and threats to validity. Section 8 reviews related work.  Section 9 concludes the paper. Finally,  Section 10 presents the details of data availability.
@@ 2.1. Issue-Commit Linking Recovery
An example of an issue, the incorrectly top-ranked commit, and the correct commit. The input issue, with summary and description, is shown below. Issue ID: NETBEANS-803 Summary: nb-javac 11 upgrade in NetBeans Description: Should cover below tasks in NetBeans - nb-javac 11 testing : Run tests for modules java.completion, java.editor, java.editor.base, java.hints, java.source, java.source.base, lib.nbjavac Update libs.javacimpl and libs.javacapi jars, upload updated nb-javac jars Upload nb-javac module jars in update center   The top-ranked commit message in the initial results is shown below but is incorrect. Commit ID: 4fd115aeae3b8423de9ced22d52914e60a1c5800. Updation for external nb-javac jar in libs.javacapi and libs.javaimpl modules with nb-javac jar for jdk-12   The correct commit message is shown below. Commit ID: 3055661e4dd1c7d587012c917cbec31b27ae9e34 Uptake nb-javac 11 jars for java tests runtime
@@ 3.1. Ground Truth Dataset Selection
We selected datasets from four recent studies, covering 20 open-source software projects over a span of 20 years. We include the same issues and true links from these datasets, resulting in a total of 9,319 issues in the benchmark. For datasets with incomplete information detected through a manual check, we re-fetched the necessary data, such as issue comments, code diffs, and committed files (i.e., the source code files after applying the commits).
@@ 5 Function EvalSetGen(test_set_true_links, commits_pool):
[…]
@@
[…]
@@
[…]
@@ 4.1. Vector Database Indexing and Retrieval
To retrieve commits from the vector database, we compute the squared L2 (Euclidean) distance between the query embedding and the stored embeddings, as we use the vector database through LangChain (Chase, 2022), where L2 distance is used by default. The computation is given by the following equation:
(3) $distance(\mathbf{q},\mathbf{e}_{i})=|\mathbf{q}-\mathbf{e}_{i}|_{2}^{2}$
where $e_{i}$ is the element of the commit message embeddings $\mathbf{E}=\langle\mathbf{e}_{1},\mathbf{e}_{2},\dots,\mathbf{e}_{n}\rangle$ that are indexed by the vector database for efficient access, and $\mathbf{q}$ is an issue embedding. A shorter distance between the query issue and the commit message indicates greater similarity, enabling the ranking of commit messages and the generation of a list of candidate commits.
@@ 4.2. LLM-Assisted Reranking
In detail, we adopt a zero-shot approach guided by prompts with ChatGPT (gpt-4o), following recent studies (Wang et al., 2024; Islam and Moushi, 2024; Fuchß et al., 2025), which demonstrate gpt-4o advanced capabilities in understanding complex textual relationships. The parameter $k$ controls the additional cost incurred for increasing precision. While it is possible to rerank all fetched commits for higher precision, this requires a longer prompt for the LLM, requiring more computation resources. For the value of $k$, we select $k$=10 for its balance of precision and efficiency. We later show that this allows a high Precision@1 without incurring a high cost. This will be discussed in Subsection 6.3.
@@ 5.3. Evaluation Metrics
[…]
@@ 6.1. RQ1: How does the state-of-the-art tool perform on a realistic evaluation dataset?
Then, we ran EALink on the evaluation dataset constructed under the Realistic Distribution Setting. Unlike the original evaluation dataset of Zhang et al. (Zhang et al., 2023b), which had a fixed number of 99 false links per issue, our evaluation generates more false links for repositories with a larger number of commits. This results in an average number of false links per issue of 1529.78. In the second column of Table 2, we present the results of running the tool on our evaluation dataset, which shows a significant performance drop compared to the original evaluation dataset. Across the 20 projects, EALink’s Precision@1 decreases to 14.43%, a decline of 55.63%. Similarly, Hit@10 drops from nearly 60% to about 30%, reducing by half. On the six projects provided by EALink (including Ambari, Calcite, and four others), the average Precision@1 decreases from 53.67% to 28.05%. Additionally, we evaluated the traditional IR method, VSM, on the same evaluation dataset. Its results are presented in Table 4. Surprisingly, EALink underperforms the traditional IR method, VSM. VSM achieves a Precision@1 of 46.59% and demonstrates better performance across all other metrics. These results suggest that the use of VSM would be preferred over EALink in a realistic setting.
@@ 6.2.1. Comparison of Effectiveness
P@1 (Hit@1) 53.67 81.59 89.58($\uparrow 66.90\%$)
P@10 8.94 11.08 11.08($\uparrow 23.93\%$)
Hit@10 72.56 92.44 92.44($\uparrow 27.39\%$)
Recall@10 69.03 89.43 89.43($\uparrow 29.55\%$)
MRR 60.13 85.57 90.80($\uparrow 51.00\%$)
NDCG@1 33.86 51.48 56.52($\uparrow 66.92\%$)
NDCG@10 41.33 56.07 58.03($\uparrow 40.41\%$)
@@
P@1 (Hit@1) 14.43 46.59 64.03 75.91 ($\uparrow 426.1\%$)
P@10 3.64 8.70 10.82 10.82 ($\uparrow 197.3\%$)
Hit@10 30.76 70.94 85.35 85.35 ($\uparrow 177.5\%$)
Recall@10 28.06 67.22 81.48 81.48 ($\uparrow 190.4\%$)
MRR 20.21 55.10 71.59 79.92 ($\uparrow 295.4\%$)
NDCG@1 9.10 29.39 40.40 47.89 ($\uparrow 426.3\%$)
NDCG@10 15.52 39.45 49.37 52.34 ($\uparrow 237.2\%$)
@@
To ensure a fair comparison, we first evaluated the vector database and EasyLink on the dataset constructed using the EALink method (Zhang et al., 2023b), which generates 99 false links per issue and includes the six projects provided by EALink. The results, shown in Table 3, indicate that EasyLink achieves a Precision@1 of 89.58%, significantly outperforming EALink’s 53.67%. The vector database method also outperforms EALink, achieving a Precision@1 of 81.59%.
Furthermore, we compared them using our more realistic evaluation dataset. As presented in Table 4, the vector database approach achieves an average Precision@1 of 64.03%, a significant improvement over EALink’s 14.43%. With LLM-assisted reranking, EasyLink further enhances Precision@1 by an average of 11.88%, reaching 75.91%, representing a 426.1% increase compared to EALink. Following the Mann-Whitney U test (Mann and Whitney, 1947), the improvement of EasyLink over EALink in every metric is statistically significant (p-value $<$ 0.01) and exhibits a large effect size (Cohen, 2013) (Cohen’s D $>$ 0.8). These results demonstrate that EasyLink is effective in issue-commit link recovery and that LLMs can be effectively leveraged to enhance performance.

### v2

@@ Abstract.
Inspired by these observations, we propose EasyLink, which utilizes a vector database as a modern Information Retrieval technique. To address the long-standing problem of the semantic gap between issues and commits, EasyLink leverages a large language model to rerank the commits retrieved from the database. Under our evaluation, EasyLink achieves an average Precision@1 of 75.03%, improving over the state-of-the-art by over four times. Additionally, this paper provides practical guidelines for advancing research in issue-commit link recovery.
@@
Jieke Shi and Ting Zhang are the corresponding authors.
journalyear: 2026copyright: rightsretainedconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3764530isbn: 979-8-4007-2025-3/26/04
@@ 1. Introduction
Prior studies (Bachmann and Bernstein, 2009; Romo et al., 2014) have revealed that many issue-commit links can be missing during the development of large-scale projects. Manually recovering these links is not only time-consuming but error-prone, even for experienced developers (Ruan et al., 2019). To address this, several studies (Zhang et al., 2023b; Lin et al., 2021; Dong et al., 2022; Ruan et al., 2019; Sun et al., 2017a; Rath et al., 2018) have proposed learning-based approaches to automatically recover issue-commit links, achieving strong performance on datasets collected from open-source repositories. The evaluation method requires the tool to distinguish fix commits (i.e., “true links”) from plausible but non-fix commits (i.e., “false links”). However, prior studies’ evaluations often lack realism. ❶ Some studies use an unrealistic time window to select commits. For example, studies (Ruan et al., 2019; Nguyen et al., 2012; Sun et al., 2017b; Bachmann et al., 2010) rely on a narrow 7-day time window to select potential fix commits, which may miss many true links. Another example involves studies (Dong et al., 2022; Rath et al., 2018) that require the issue close time to select commits, but in practice, when commits are missing, the issue may not have a close time. ❷ Another limitation is the unrealistic false link distributions in prior evaluation datasets. Some (Lin et al., 2021; Ruan et al., 2019) use evaluation datasets where the number of false and true links is equal, an unrealistic assumption since false links far outnumber true links in reality (Dong et al., 2022). Recently, Zhang et al. (Zhang et al., 2023b) addressed this issue by using an imbalanced dataset with a fixed number of false links per issue. Still, this method overlooks a crucial factor: a higher commit frequency results in a larger pool of plausible commits for the tool to differentiate from the actual fix commit, making the task inherently more difficult.
@@
In this paper, we propose EasyLink, a novel method that integrates modern IR techniques and Large Language Models (LLMs) to enhance issue-commit linking performance in realistic settings by capturing deeper semantic relationships between issues and commits. EasyLink operates in two stages. The first stage fetches a set of commits that share similarities to a given issue, and the second stage reranks them by their relevance to the issue. Concretely, in the first stage, EasyLink leverages recent advances in Information Retrieval by using a modern vector database (optimized for high-dimensional similarity search) (Douze et al., 2024) to fetch the most similar commits for each issue. In the second stage, EasyLink prompts an LLM (GPT-4o (OpenAI, 2024)) to rerank the results. We use LLMs for their strong ability to capture the semantic relationship between issues and commits, bridging the semantic gap, as the most similar commit does not guarantee that it is the fix for the issue (Zhang et al., 2023b; Lin et al., 2021; Dong et al., 2022). EasyLink achieves an average Precision@1 of 75.03%, outperforming EALink (Zhang et al., 2023b) by four times in our realistic datasets and by more than 30% in Precision@1 on EALink’s original evaluation setup. We also conduct an ablation analysis on each of EasyLink’s two stages. Our results demonstrate that advances in IR, such as context-aware dense embeddings (Wang et al., 2020; Song et al., 2020; Neelakantan et al., 2022) and efficient vector search (Douze et al., 2024; Wang et al., 2021), have been largely overlooked in the software traceability literature. Notably, even the out-of-the-box use of a vector database achieves a high average Precision@1 of 61.57%. Additionally, the strong performance of the reranking step, which improves Precision@1 by 13.46%, highlights the capability of LLMs to bridge the semantic gap between issues and commits (Rath et al., 2018; Guo et al., 2017). Finally, we discuss the lessons learned from our work, such as the need to consider modern IR baselines, for future research on software traceability. Our implementation has been made available at (Package, 2025).
@@
    A new state-of-the-art for issue-commit linking: We propose EasyLink, which leverages modern IR techniques, including an off-the-shelf database FAISS (Douze et al., 2024), and addresses the problem of semantic gap (Rath et al., 2018; Guo et al., 2017) by prompting an LLM to rerank the retrieved results. In our evaluation, EasyLink improves over EALink in average Precision@1 from 14.43% to 75.03%.
The rest of the paper is organized as follows. Section 2 outlines the background of this work, including the limitations of existing work. Section 3 details the evaluation dataset construction under the Realistic Distribution Setting (RDS). Section 4 introduces EasyLinkSection 5 describes the experimental setup. Section 6 presents the experimental results. Section 7 revisits the classical baseline, discusses the lessons learned and threats to validity.  Section 8 reviews related work. Finally,  Section 9 concludes the paper.
@@ 2.1. Issue-Commit Linking Recovery
An example of an issue, the incorrectly top-ranked commit, and the correct commit. The input issue, with summary and description, is shown below.
Issue ID: NETBEANS-803
Summary:
nb-javac 11 upgrade in NetBeans
Description:
Should cover below tasks in NetBeans - nb-javac 11 testing :
• Run tests for modules java.completion, java.editor,
 java.editor.base, java.hints, java.source,
 java.source.base, lib.nbjavac
• Update libs.javacimpl and libs.javacapi jars, upload updated
 nb-javac jars
• Upload nb-javac module jars in update center
  The top-ranked commit message in the initial results is shown below but is incorrect.
Commit ID: 4fd115aeae3b8423de9ced22d52914e60a1c5800.
Updation for external nb-javac jar in libs.javacapi and
libs.javaimpl modules with nb-javac jar for jdk-12
  The correct commit message is shown below.
Commit ID: 3055661e4dd1c7d587012c917cbec31b27ae9e34
Uptake nb-javac 11 jars for java tests runtime
@@ 3.1. Ground Truth Dataset Selection
We selected datasets from four recent studies, covering 20 open-source software projects over a span of 20 years. These datasets use explicit issue tags (e.g., “#123”, “JIRA-456”) in commit messages to construct the true links. We include the same issues and true links from these datasets, resulting in a total of 9,319 issues in the benchmark. For datasets with incomplete information detected through a manual check, we re-fetched the necessary data, such as issue comments, code diffs, and committed files (i.e., the source code files after applying the commits).
@@ 5 Function EvalSetGen(test_set_true_links, commits_pool):
13    candidate_commits $\leftarrow$ { commit $\in$ commits_pool $\mid$
@@
18    false_links $\leftarrow$ {(issue, commit, label=0) $\mid$
@@
$\epsilon$ $=\text{one year.}$
Algorithm 1 details the construction of the evaluation dataset. We first split the entire ground-truth dataset into training and test sets following a 4:1 ratio (Zhang et al., 2023b; Dong et al., 2022; Ruan et al., 2019). The inputs to the algorithm are the true links from the test set and the commits pool for the project, which is prepared using the process described in Subsection 3.2. First, following the method used by EALink (Zhang et al., 2023b), we randomly sample up to 1,000 unique issue IDs from the test set, forming the list $\mathit{sampled\_issue\_id\_list}$. If fewer than 1,000 unique issues are available, all issues are included. Then, for each $\mathit{issue\_id}$ in this list, we extract the true links for that issue from the test set. We use the term “links” (plural) because an issue may be linked to more than one commit (Zhang et al., 2023b). In line 6 of the algorithm, we extract only the issue information for the issue being processed (including issue ID, summary, description, etc.), denoted as $\mathit{issue}$. Next, as shown in lines 7 to 10 of the algorithm, we filter the $\mathit{candidate\_commits}$ to include all commits whose creation time is no earlier than the issue’s creation time and strictly earlier than one year after the issue’s creation time. Additionally, the commit should not be the fix commit of the issue. We then construct the $\mathit{false\_links}$ by pairing the $\mathit{issue}$ with each commit in $\mathit{candidate\_commits}$ and labeling the pair as 0. Finally, we append both the $\mathit{true\_links}$ and the $\mathit{false\_links}$ for the issue to the evaluation dataset and then proceed to process the next issue. Note that this algorithm is executed separately for each project, as each project has its own test dataset and commits pool. Following prior work (Zhang et al., 2023b; Ruan et al., 2019), issue tags (e.g., “#123”, “JIRA-456”) in the issue and commit text are removed to prevent potential data leakage.
@@ 4.1. Vector Database Indexing and Retrieval
To retrieve commits from the vector database, we compute the cosine similarity score, following prior studies (Zhang et al., 2023b; Ruan et al., 2019; Rath et al., 2018). The computation is given by the following equation:
(3) $\operatorname{sim}(\mathbf{q},\mathbf{e}_{i})\;=\;\frac{\mathbf{q}\cdot\mathbf{e}_{i}}{\lVert\mathbf{q}\rVert_{2}\,\lVert\mathbf{e}_{i}\rVert_{2}}$
where $\mathbf{e}_{i}$ is the $i$-th element of the commit message embeddings $\mathbf{E}=\langle\mathbf{e}_{1},\mathbf{e}_{2},\dots,\mathbf{e}_{n}\rangle$ that are indexed by the vector database for efficient access, and $\mathbf{q}$ is the issue embedding. A higher score between the query issue and the commit message indicates greater similarity, enabling the ranking of commit messages and the generation of a list of candidate commits.
@@ 4.2. LLM-Assisted Reranking
In detail, we adopt a zero-shot approach guided by prompts with ChatGPT (gpt-4o), following recent studies (Wang et al., 2024; Islam and Moushi, 2024; Fuchß et al., 2025), which demonstrate gpt-4o advanced capabilities in understanding complex textual relationships. The parameter $k$ controls the additional cost incurred for increasing precision. While it is possible to rerank all fetched commits for higher precision, this requires a longer prompt for the LLM, requiring more computation resources. For the value of $k$, we select kk=10 for its balance of precision and efficiency. We later show that this allows a high Precision@1 without incurring a high cost. This will be discussed in Subsection 6.3.
@@ 5.3. Evaluation Metrics
    (6) $\text{Recall@k}=\frac{1}{|Q|}\sum_{i\in Q}\frac{\text{Rel}_{i}}{\text{TotalRel}_{i}},$
@@ 6.1. RQ1: How does the state-of-the-art tool perform on a realistic evaluation dataset?
Then, we ran EALink on the evaluation dataset constructed under the Realistic Distribution Setting. Unlike the original evaluation dataset of Zhang et al. (Zhang et al., 2023b), which had a fixed number of 99 false links per issue, our evaluation generates more false links for repositories with a larger number of commits. This results in an average number of false links per issue of 1,530. In the second column of Table 2, we present the results of running the tool on our evaluation dataset, which shows a significant performance drop compared to the original evaluation dataset. Across the 20 projects, EALink’s Precision@1 decreases to 14.43%, a decline of 55.63%. Similarly, Hit@10 drops from nearly 60% to about 30%, reducing by half. On the six projects provided by EALink (including Ambari, Calcite, and four others), the average Precision@1 decreases from 53.67% to 28.05%. Additionally, we evaluated the traditional IR method, VSM, on the same evaluation dataset. Its results are presented in Table 4. Surprisingly, EALink underperforms the traditional IR method, VSM. VSM achieves a Precision@1 of 46.59% and demonstrates better performance across all other metrics. These results suggest that the use of VSM would be preferred over EALink in a realistic setting.
@@ 6.2.1. Comparison of Effectiveness
P@1 (Hit@1) 53.67 81.83 90.04($\uparrow 67.77\%$)
P@10 8.94 11.14 11.14($\uparrow 24.61\%$)
Hit@10 72.56 92.89 92.89($\uparrow 28.02\%$)
Recall@10 69.03 89.93 89.93($\uparrow 30.28\%$)
MRR 60.13 85.87 91.38($\uparrow 51.97\%$)
NDCG@1 33.86 51.63 56.81($\uparrow 67.78\%$)
NDCG@10 41.33 56.31 58.07($\uparrow 40.50\%$)
@@
P@1 (Hit@1) 14.43 46.59 61.57 75.03 ($\uparrow 420.0\%$)
P@10 3.64 8.70 10.09 10.09 ($\uparrow 177.2\%$)
Hit@10 30.76 70.94 83.38 83.38 ($\uparrow 171.1\%$)
Recall@10 28.06 67.22 78.84 78.84 ($\uparrow 181.0\%$)
MRR 20.21 55.10 69.21 78.52 ($\uparrow 288.5\%$)
NDCG@1 9.10 29.39 38.85 46.86 ($\uparrow 415.0\%$)
NDCG@10 15.52 39.45 47.98 51.08 ($\uparrow 229.1\%$)
@@
To ensure a fair comparison, we first evaluated the vector database and EasyLink on the dataset constructed using the EALink method (Zhang et al., 2023b), which generates 99 false links per issue and includes the six projects provided by EALink. The results, shown in Table 3, indicate that EasyLink achieves a Precision@1 of 90.04%, significantly outperforming EALink’s 53.67%. The vector database method also outperforms EALink, achieving a Precision@1 of 81.83%.
Furthermore, we compared them using our more realistic evaluation dataset. As presented in Table 4, the vector database approach achieves an average Precision@1 of 61.57%, a significant improvement over EALink’s 14.43%. With LLM-assisted reranking, EasyLink further enhances Precision@1 by an average of 13.46%, reaching 75.03%, representing a 420.0% increase compared to EALink. Following the Mann-Whitney U test (Mann and Whitney, 1947), the improvement of EasyLink over EALink in every metric is statistically significant (p-value << 0.01) and exhibits a large effect size (Cohen, 2013) (Cohen’s D >> 0.8). These results demonstrate that EasyLink is effective in issue-commit link recovery and that LLMs can be effectively leveraged to enhance performance.